From 200gm of co2,10^21 molecules are removed. How many grams of co2 are left?

Grade:11

1 Answers

Meghana C S
5 years ago
Number of moles of CO2 in 200gm = 200÷44=4.54 moles 
No: of molecules in 4.54 moles is 4.54 ×6.022 ×10^23
When 10^21 removed remaining
molecules=2733×10^21
Converting back to moles (÷by avogadro no:)=453×10^-2 
Therefore, mass of co2 left =453×10^-2×44=199.32gm